find the coordinates of the point L that divides segment MN into a ratio of ML:LN to 1:3; M(-4,0) and N(0,4). Write your answer in coordinate form (x,y)


1
Expert's answer
2021-10-11T12:27:04-0400

Let the coordinates of point "L" be "(x,y)."

Given that "ML:LN" is "1:3" .

Using section formula, we get:

"(x,y)=(\frac{-12+0}{4},\frac{0+4}{3})\\
\therefore x=-3,y=\frac{4}{3}"

So, coordinates of point "L" be "(-3,\frac{4}{3})."